OpenStudy (anonymous):

FREE MEDALS!! IF YOU ANSWER! Match the vertices of each triangle with its classification: here are the Vertices: 1. (3, 5), (5, 1), (3, 1) 2. (2, 3), (0, 0), (4, 0) 3. (1, 2), (3, 2), (6, 5) 4. (3, 3), (3, 6), (0, 3) and here are the classifications: A) acute isosceles B) right scalene C) right isosceles D) obtuse scalene.

OpenStudy (anonymous):

1/ the first one we notice A(3;5) & C (3,1) have the same x => vertical line and B & C the same y value => horizontal. Right Triangle

OpenStudy (anonymous):

what about the other ones?

OpenStudy (anonymous):

2/ acute isoscele 3/ obtuse scalene. 4/ right isosceles

OpenStudy (anonymous):

sketching the points on a grid will give you ideas. 2/ B & C make horizontal line and A is in the middle of the BC distance (projection on x) 4/ (3, 3), (3, 6), (0, 3), A & B vertical line; B&C horizontal => right angle and AB=AC (AB=sqrt((3-3)²+(3-6)²) and AC=sqrt((3-0)²+(3-3)²))

OpenStudy (anonymous):

THANK YOU SOOOOOOOO MUCH!!!

OpenStudy (anonymous):

I get it now!!! (: